OpenStudy (anonymous):

Please help:Solve. (2x^2) - 8x -14 = 0 1) x= 4± 3 √2 2) x= 4± √3

OpenStudy (anonymous):

\[x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

I tried doing that, i got down to 4±√44 / 2

OpenStudy (anonymous):

\[x = \frac{ 8 \pm \sqrt{64 + 14 \times 4}}{4}\]

OpenStudy (anonymous):

try it now

OpenStudy (anonymous):

8±√78 ?

OpenStudy (anonymous):

8±√120/4